-
#200
by
psloss
on 18 Dec, 2007 14:36
-
Backup pump is online already; should be getting to stable replenish on the LO2 side soon...
-
#201
by
psloss
on 18 Dec, 2007 14:38
-
Now in stable replenish on both sides (controlled switching). The red teams will be headed to the pad shortly.
-
#202
by
Chris Bergin
on 18 Dec, 2007 14:49
-
-
#203
by
Chris Bergin
on 18 Dec, 2007 14:53
-
And up in the elevator.
"*Bing* Menswear"
-
#204
by
Chris Bergin
on 18 Dec, 2007 15:00
-
-
#205
by
Chris Bergin
on 18 Dec, 2007 15:08
-
-
#206
by
Chris Bergin
on 18 Dec, 2007 15:09
-
-
#207
by
Chris Bergin
on 18 Dec, 2007 15:13
-
Voltage readings confirm console readouts on the health of the four sensors/circuits.
-
#208
by
Chris Bergin
on 18 Dec, 2007 15:17
-
TDR Engineer's laptop keeps freezing, I've noticed. Has had to reboot it twice already.
-
#209
by
Chris Bergin
on 18 Dec, 2007 15:21
-
-
#210
by
triddirt
on 18 Dec, 2007 15:21
-
Chris Bergin - 18/12/2007 11:17 AM
TDR Engineer's laptop keeps freezing, I've noticed. Has had to reboot it twice already.
"Keeps Freezing" No pun intended I'm sure :cool:
-
#211
by
Chris Bergin
on 18 Dec, 2007 15:29
-
triddirt - 18/12/2007 4:21 PM
Chris Bergin - 18/12/2007 11:17 AM
TDR Engineer's laptop keeps freezing, I've noticed. Has had to reboot it twice already.
"Keeps Freezing" No pun intended I'm sure :cool:
I'm available for children's parties and office functions. Plenty more cryo related puns in my back pocket

TDR checks on sensor 1 and 2. Any body lanuage readers on here (I am not being serious!)
-
#212
by
Chris Bergin
on 18 Dec, 2007 15:31
-
And now reading data from LH2 ECO sensor #3
-
#213
by
Chris Bergin
on 18 Dec, 2007 15:34
-
Maybe they've found something! (Guessing)
-
#214
by
Chris Bergin
on 18 Dec, 2007 15:41
-
And now TDR'ing LH2 ECO #4....and through to the five percent sensors.
-
#215
by
Chris Bergin
on 18 Dec, 2007 15:52
-
Inspection team at MLP level now.
-
#216
by
Chris Bergin
on 18 Dec, 2007 16:00
-
TDR testing phase 1 complete on all LH2 ECO sensors (low level/5 percent).
-
#217
by
Chris Bergin
on 18 Dec, 2007 16:13
-
TDR Troubleshooting (manually inside the MLP) is complete.
-
#218
by
Chris Bergin
on 18 Dec, 2007 16:14
-
TDR team now leaving the MLP. TDR will remain active, as console operators will watch the sensors during detanking.
-
#219
by
Chris Bergin
on 18 Dec, 2007 16:24
-
L2 sources note initial TDR data point to the problem being with the LH2 Feedthrough Connector.
This - at this stage - would confirm the engineering thought process on this prior to the tanking test:
http://www.nasaspaceflight.com/content/?cid=5309This could get interesting, as extensive MSFC testing, gained via presentation slides on L2 yesterday, showed they could not find a problem with the connector during lab testing.